OpenStudy (anonymous):

I am Almost done with my assignment could someone please help with this ?? 9th grade algebra

OpenStudy (anonymous):

A store had 235 MP3 players in the month of January. Every month, 30% of the MP3 players were sold and 50 new MP3 players were stocked in the store. Which recursive function best represents the number of MP3 players in the store f(n) after n months? f(n) = 0.7 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 237 - 0.7 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 0.3 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 237 + 0.7 × f(n - 1) + 50, f(0) = 235, n > 0

OpenStudy (anonymous):

@douglaswinslowcooper Could you help me out again ?

OpenStudy (anonymous):

i'LL TRY. f(1) = 235 starting out in January month 1 f(2) = (0.70)(235) + 50 70% remained unsold and 50 were added f(2) = (0.70) f(1) + 50 what will f(3) be, using tis same logic?

OpenStudy (anonymous):

I see they want January to be n=0, so f(0)= 235 f(1) = (0.7) f(0) + 50 f(2) = (0.7) f(1) + 50 next?

OpenStudy (anonymous):

so since 0.7 is being multiplied to f(n) then C would be out right ?

OpenStudy (anonymous):

Yes. Good start. What is f(3) following my pattern?

OpenStudy (anonymous):

(0.7) f(2) + 50

OpenStudy (anonymous):

Wouldn't the answer come out as A since you are multiplying 0.7

OpenStudy (anonymous):

"note" not "not"!

OpenStudy (anonymous):

f(n) goes toward 0.3 f(n) = sold = 50 added f(n) = 50/0.3 = 167 as n gets >> 1

OpenStudy (anonymous):

Glad to have been helpful.

OpenStudy (anonymous):

Thank yo very much tho !